OpenStudy (debpriya):

a number N when divided by a divisor D leaves a remainder of 19. When 3N is divided by D, the remainder is 14. What is the value of D?(a.23 b.37 c.43 d.47)?

OpenStudy (vishweshshrimali5):

We will use this formula: Dividend = Divisor * Quotient + Remainder

OpenStudy (vishweshshrimali5):

So, in first case... We have: \[N = D* Q_1 + 19\]

OpenStudy (vishweshshrimali5):

Any doubt so far?

OpenStudy (vishweshshrimali5):

Similarly for second case: \[3N = D*Q_2 + 14\]

OpenStudy (vishweshshrimali5):

@debpriya any doubts so far?

OpenStudy (debpriya):

nope got it :) ...what after that?

OpenStudy (vishweshshrimali5):

\[\implies 3*D*Q_1 + 57 = D*Q_2 + 14\] \[\implies D*(Q_2 - 3*Q_1) = 43\]

OpenStudy (vishweshshrimali5):

Now can you guess what we can do after this?

OpenStudy (vishweshshrimali5):

Hint: 43 is a prime number...

OpenStudy (kindaljoshconall):

hi need help

OpenStudy (vishweshshrimali5):

@kindaljoshconall can you tag me in the question you want help in?

OpenStudy (kindaljoshconall):

sure

OpenStudy (kindaljoshconall):

its on science

OpenStudy (vishweshshrimali5):

I can try :)

OpenStudy (knov):

doesn't matter in what section your question is. you can tag anyone in any section.

OpenStudy (kindaljoshconall):

kk ty i fanned u

OpenStudy (vishweshshrimali5):

@debpriya any guesses?

OpenStudy (debpriya):

none so far :(

OpenStudy (vishweshshrimali5):

No problem, see, 43 is a prime number... thus it will have only two factors 1 and 43... Also, we have from the equation that D*(Q2 - 3*Q1) = 43 that means D and (Q2 - 3*Q1) are also the factors of 43... Now, there arise two cases... D = 1 or D = 43...Now, since every number is a multiple of itself, dividing N by 1 should not leave any remainder... Thus, only possibility is D = 43.

OpenStudy (vishweshshrimali5):

If you have any kind of query regarding this solution, please feel free to ask

OpenStudy (debpriya):

oh I got it...thank you so much for your help :)

OpenStudy (vishweshshrimali5):

:) Glad to be of help

OpenStudy (kindaljoshconall):

hes a morning helper lol